A U.S. manufacturing company operating a subsidiary in an LDC (less-developed country) shows the following results:

U.S. LDC
Sales (units) 100,000 20,000
Labor (hours) 20,000 15,000
Raw materials (currency) $ 20,000 FC* 20,000
Capital equipment (hours) 60,000 5,000

*Foreign Currency unit

a. Calculate partial labor and capital productivity figures for the parent and subsidiary. (Round your answers to 2 decimal places.)

U.S. LDC
Labor productivity
Capital productivity

b. Compute the multifactor productivity figures for labor and capital together. (Round your answers to 2 decimal places.)

U.S. LDC
Multifactor productivity

c. Calculate raw material productivity figures (units/$ where $1 = FC 10). (Round your answers to 2 decimal places.)

U.S. LDC
Raw material productivity
Answer #1

a)

Labor productivity = Sales/Labor hours

US = 100,000/20,000 = 5 units per labor hour

LDC = 20,000/15,000 = 1.33 units per labor hour

Capital productivity = Sales/Capital hours

US = 100,000/60,000 = 1.67 units per capital hour

LDC = 20,000/5,000 = 4 units per capital hour

b)

Multifactor productivity = Sales/(Labor + Capital)

US = 100,000/(20,000+60,000) = 1.25 units per hour

LDC = 20,000/(15,000+5,000) = 1 unit per hour

c)

Raw material productivity = Sales/Raw material

US = 100,000/20,000 = 5 units per dollar

LDC = 20,000/20,000 = 1 unit per FC or 1 unit per 0.1 dollar or 10 units per dollar